# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ VII Concluding Remarks
continuous-time population dynamics Stochastic approximation theory Coordinating Anticoordinating Mixture of both $t\ \rightarrow\ \infty$EquilibratesEquilibratesEquilibratesThe fluctuations vanish Population EquilibratesEquilibratesMay Fluctuate Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both $t\ \rightarrow\ \infty$$t\ \rightarrow\ \infty$ Population Population
@@
Figure 2: [Revisit]-The connection between the asymptotic behavior of finite populations of interacting agents evolving based on (3) and their associated continuous-time population dynamics (7) based on Stochastic approximation theory. As population size approaches infinity, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players vanishes with probability $1$.
@@
[…]
A Caratheodory solution of (4) on $[0,t_{1}]\subset\mathbb{R}_{\geq 0}$ is an absolutely continuous map $\bm{x}^{o}:[0,t_{1}]\rightarrow\mathbb{R}^{n}$ such that for almost all $t\in[0,t_{1}]$, $\dot{\bm{x}}^{o}(t)\in\bm{\mathcal{V}}(\bm{x}^{o}(t))$ [51]. From now on, by a solution of (4), we mean its Caratheodory solution. Robotic manipulators interacting with objects, high-speed supercavitation vehicles, and power systems can be well modeled and analyzed within the framework of differential inclusions [51, 52, 53]. A point $\bm{x}^{*}\in\mathbb{R}^{n}$ that satisfies ${\mathbf{0}}\in\bm{\mathcal{V}}(\bm{x}^{*})$ is an equilibrium of (4) [51]. The set-valued dynamical system induced by the differential inclusion (4) is denoted by $\bm{\Phi}(\bm{x}_{0})$ and defined as the set of all solutions of (4) with initial condition $\bm{x}_{0}$ [38].
Example A 1.
Consider the scalar differential inclusion (4) with $x\in\mathbb{R}$ and
$\mathcal{V}\big{(}x(t)\big{)}=\begin{cases}\{-1-x(t)\}&\text{if }x(t)<0,\\ [-1,1]&\text{if }x(t)=0,\\ \{1-x(t)\}&\text{if }x(t)>0.\end{cases}$
For any positive initial condition $x(0)$, the solution of (4) equals $x(t)=1+e^{-t}(x(0)-1)$. For any negative initial condition $x(0)$, the solution of (4) equals $x(t)=-1+e^{-t}(1+x(0))$. For the initial condition $x(0)=0$, however, the solution $x(t)$ is not unique, see Figure 3. The set valued dynamical system $\bm{\Phi}$ for initial condition $x(0)=0$ contains $\{0,e^{-t}-1,1-e^{-t}\}$, and for other initial conditions will be a singleton. As for equilibrium points of (4), we need to find the points $x^{*}$ at which $0\in\mathcal{V}(x^{*})$, yielding $-1$, 00, and $1$.
Figure 3: 1. The upper panel represents the set-valued map $\mathcal{V}(x)$. The lower panel represents three solutions of $\dot{x}(t)\in\mathcal{V}(x)$ with 00 as the initial state, i.e., $0,e^{-t}-1,$ and $1-e^{-t}$.
A set-valued map $\bm{\mathcal{V}}:\bm{\mathcal{X}}\rightrightarrows\bm{\mathcal{Y}}$ is called upper semicontinuous if for every sequence $\langle\bm{x}_{n}\rangle_{{n=0}}^{{\infty}}$, $\bm{x}_{n}\in\bm{\mathcal{X}}$, with limit point $\bm{x}_{0}\in\bm{\mathcal{X}}$ and every sequence $\langle\bm{y}_{n}\rangle_{{n=0}}^{{\infty}}$, $\bm{y}_{n}\in\bm{\mathcal{V}}(\bm{x}_{n})$, with limit point $\bm{y}_{0}$, we have $\bm{y}_{0}\in\bm{\mathcal{V}}(\bm{x}_{0})$. When an upper semicontinuous map is equipped with three additional properties of nonemptiness, convexity, and boundedness, the map is good upper semicontinuous. As for the following scalar set-valued map
$\mathcal{V}(x(t))=\begin{cases}\{-1-x(t)\}&\text{if }x(t)<0,\\ [0,1]&\text{if }x(t)=0,\\ \{1-x(t)\}&\text{if }x(t)>0,\end{cases}$
consider the sequence $\langle x_{n}\rangle_{n}$, where $x_{n}=-0.5/(n+1)$, for $n=0,1,\ldots$. The sequence is associated with $\langle y_{n}\rangle_{n}$, where $y_{n}=-1+0.5/(n+1)$. The limit point of the sequence $\langle x_{n}\rangle_{n}$ is $x_{0}=0$ and that of the sequence $\langle y_{n}\rangle_{n}$ is $y_{0}=-1$. However, $y_{0}\notin\mathcal{V}(x_{0})$. Hence, the map is not upper semicontinuous, because it is not so at $x=0$. A sequence $\langle\epsilon_{k}\rangle_{{k=0}}^{{\infty}}$ of positive constants that converges to zero is called vanishing.
@@
[38, Theorem 3.5 and Corollary 3.9] For a vanishing sequence of $\epsilon>0$, let $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}\rangle_{\epsilon>0}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k}$ is a Markov chain and let ${\mu}^{\epsilon}$ be an invariant probability measure of $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{{}_{k=0}^{\infty}}}$. Let ${\mu}$ be a limit point of $\langle{\mu}^{\epsilon}\rangle_{\epsilon>0}$ in the topology of weak convergence. Then the support of ${\mu}$ is contained in $\mathrm{BC}(\bm{\Phi})$, where $\bm{\Phi}$ is the dynamical system induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
@@ VII-A Proof of Theorem 1
Assume by contradiction that $\lim_{k\to\infty}{{\mu}}^{\epsilon_{k}}(\bm{\mathcal{{O}}})\not\to 1$. This implies that for some positive constant $\delta<1$, there exists a subsequence $\langle\alpha_{k}\rangle_{k=1}^{\infty}=\alpha_{1},\alpha_{2},\ldots$ of $\langle\epsilon_{k}\rangle_{k=0}^{\infty}$ such that ${{\mu}}^{\alpha_{k}}(\bm{\mathcal{{O}}})<\delta$ for each $k$. As the state space $\bm{\mathcal{X}}_{0}$ is compact, the family of probability measures defined on the space $\bm{\mathcal{X}}_{0}$ is tight [50]. Then, in view of Prokhorov’s theorem, every sequence of the probability measures, including $\langle{\mu}^{\alpha}\rangle$ has a weak limit point, which is a probability measure ${\mu^{\prime}}$. This implies that there exists a subsequence $\langle\beta_{k}\rangle_{k=1}^{\infty}=\beta_{1},\beta_{2},\ldots$ of $\langle\alpha_{k}\rangle_{k=1}^{\infty}$ such that $\lim\inf_{k\to\infty}{{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})\geq{{\mu^{\prime}}}(\bm{\mathcal{{O}}})$. On the other hand, in view of A1, the support of every limit point is contained in the Birkhoff center $\mathrm{BC}(\bm{\Phi})$, i.e., ${\mu}^{\prime}(\bm{\mathcal{O}})=1$. This means that there exists some $k^{*}$ such that for all $k>k^{*}$ ${{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})>\delta$, a contradiction. ∎
@@ VII-C Proof of Lemma 1
Under 1, let $d$ denote the minimum distance between thresholds, i.e., $d=\min_{i,j\in[\mathsf{p}+\mathsf{p}^{\prime}]}|\tau_{i}-\tau_{j}|.$ Then, for any population size $\mathsf{N}$ satisfying $\mathsf{N}d>1$ and any $\bm{x}\in\bm{\mathcal{X}}_{ss}$, the sum $\bm{1}^{\top}\bm{x}$ belongs to at most one interval of the form $[\tau_{i},\tau_{i}+\frac{1}{\mathsf{N}})$ with $i\in[\mathsf{p}+\mathsf{p}^{\prime}]$.
@@
The proof of Lemma 1 is inspired by the steps taken in [38, Example 4.1].
@@
By considering the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence, it suffices to show that there exist $\langle\bm{\mathcal{V}}^{\frac{1}{\mathsf{N}}}\rangle_{{\frac{1}{\mathsf{N}}}}$ and $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ such that the conditions in Definition 4 are met. The first condition of Definition 4 can be easily verified. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit $t$ of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$. Since $k=\mathsf{N}t$, there are $\mathsf{N}$ transitions per unit time resulting in $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, which is the expected increment in the subpopulation $p$, equals the sum of the multiplication of each possible change in the subpopulation $p$ with its probability. In view of (6), there are two possible changes, i.e., $\frac{1}{\mathsf{N}}$ step increment or $\frac{1}{\mathsf{N}}$ step reduction, which results in ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\mathsf{N}\big{(}\frac{1}{\mathsf{N}}(\rho_{p}-x^{\mathsf{N}}_{p})(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-\frac{1}{\mathsf{N}}(x^{\mathsf{N}}_{p}(s^{*}(p,\bm{x}^{\mathsf{N}},{1})-1))\big{)}$ and consequently, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\rho_{p}(2-s^{*}(p,\bm{x}^{\mathsf{N}},{2}))-x^{\mathsf{N}}_{p}{+x^{\mathsf{N}}_{p}\left(s^{*}(p,\bm{x}^{\mathsf{N}},2)-s^{*}(p,\bm{x}^{\mathsf{N}},1)\right)}$. The second condition of Definition 4 is satisfied by $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}=$ $\smash{\mathsf{N}(\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}])}$ resulting in $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-{\frac{1}{\mathsf{N}}}\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}={\frac{1}{\mathsf{N}}}\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$. Now, we investigate Condition 3. To show that Condition 3 is satisfied, for a given $\delta$, we need to provide an $\epsilon_{0}$ satisfying eq. 5. We propose an $\epsilon_{0}$ which satisfies $\epsilon_{0}<\min\{d,\delta\}.$ Then, given A1, for any $\bm{x}\in\bm{\mathcal{X}}_{ss}$ there exists at most one threshold, say $\tau_{l}$, $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, that satisfies $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\frac{1}{\mathsf{N}})$, with $\frac{1}{\mathsf{N}}\leq\epsilon_{0}.$ In what follows, we show that Condition 3 holds for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ that satisfy $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$; the other values of $\bm{x}$ can be handled similarly. Let $\varepsilon_{x}=\bm{1}^{\top}\bm{x}-\tau_{l}$. Then given $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$, we have $\varepsilon_{x}<\epsilon_{0}$ and, in turn, $\varepsilon_{x}<\delta.$ Consider the point $\bm{y}$ whose all elements, except some element $j$, are equal to the corresponding elements of $\bm{x}$, and the $j^{\text{th}}$ element, $y_{j}$, equals $x_{j}-\varepsilon_{x}$. The term $|\bm{x}-\bm{y}|$ then equals $\varepsilon_{x}$, and hence the point $\bm{y}$, which satisfies $\bm{1}^{\top}\bm{y}=\tau_{l}$, belongs to the $\delta$-neighborhood of $\bm{x}$, i.e., $|\bm{x}-\bm{y}|<\delta$. For all $p\neq l$, we then have $\mathcal{V}_{p}({\bm{y}})=2(\rho_{p}-s^{*}(p,{\bm{y}},2))-{y}_{p}$ and for $p=l$, $\mathcal{V}_{p}=[-{y}_{p},\rho_{p}-{y}_{p}]$. On the other hand, the relation $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$ implies that ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ for $p=l$ equals $\rho_{l}-2x^{\frac{1}{\mathsf{N}}}_{l}$ (resp. 00) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation. For $p\neq l$, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ equals $2(\rho_{p}-s^{*}(p,\bm{x},2))-x_{p}$, and it can be easily shown that it is also equal to the unique member of $\mathcal{V}_{p}(\bm{x})$ defined in (7). By $\bm{z}=\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ and choosing $\bm{v}$ to be a member of $\bm{\mathcal{V}}({\bm{y}})$ with $v_{l}({\bm{y}})=\rho_{l}-2{y}_{l}$ (resp. $v_{l}({\bm{y}})=0$) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation, the term $|\bm{z}-\bm{v}|$ equals $|\bm{v}({\bm{y}})-\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})|=|{\bm{y}}-\bm{x}|$ which is less than $\delta$. This implies the satisfaction $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta$ and, in turn, Condition 3. As for the last condition, in view of Example 4.1 in [38], it can be shown that $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ is a martingale difference sequence. On the other hand, for all $k$, $|\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}|\leq\sqrt{\Sigma_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}(1+\rho_{l})^{2}}$ and hence $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}$ is uniformly bounded. Therefore, the conditions asserted in [38, Proposition 2.3] are satisfied and the satisfaction of the last condition is guaranteed. ∎

### v4

@@ Acknowledgment
continuous-time population dynamics Stochastic approximation theory Coordinating Anticoordinating Mixture of both tt\ $\rightarrow\ \inftyEquilibratesEquilibratesEquilibratesThe$ fluctuations vanish Population EquilibratesEquilibratesMay Fluctuate Discrete population of finite size $\mathsf{N}$ Coordinating AnticoordinatingMixture of both Discrete population $\mathsf{N}\rightarrow\infty$ Coordinating AnticoordinatingMixture of both tt\ $\rightarrow\ \inftytt\ \rightarrow\ \infty$ Population Population
@@
Figure 2: [Revisit]-The connection between the asymptotic behavior of finite populations of interacting agents evolving based on (3) and their associated continuous-time population dynamics (7) based on the stochastic approximation theory. As population size approaches infinity, the amplitude of the fluctuations in the population proportion of strategy-$\mathtt{A}$ players vanishes with probability 11.
@@
The following are mainly adopted from [klappenecker2018markov]. A discrete-time stochastic process $\langle\mathbf{X}_{k}\rangle_{k=0}^{\infty}$ is a sequence of random variables indexed by $k\in\mathbb{Z}_{\geq 0}$, where $\mathbf{X}_{k}$ is the state of the stochastic process. A Markov chain is a discrete-time stochastic process $\langle\mathbf{X}_{k}\rangle_{k=0}^{\infty}$ where $\mathbf{X}_{k+1}$ is independent of the states $\mathbf{X}_{k-1},\ldots,\mathbf{X}_{0}$ given $\mathbf{X}_{k}$. A Markov chain is finite if the space over which the states are defined, denoted by $\bm{\mathcal{S}}$, is finite. The transition probabilities of a Markov chain is denoted by $\text{Pr}_{\bm{x},\bm{y}}$ which represents the probability of reaching $\bm{y}$ at index $k+1$ given that the state at index $k$ is equal to $\bm{x}$, i.e., $\text{Pr}_{\bm{x},\bm{y}}=\mathbb{P}[\mathbf{X}_{k+1}=\bm{y}|\mathbf{X}_{k}=\bm{x}]$. The transition matrix $\mathbf{P}=[\text{Pr}_{\bm{x},\bm{y}}]$ is defined accordingly. A Markov chain is homogeneous if its transition probabilities are time independent. The invariant probability measure for a Markov chain with transition matrix $\mathbf{P}$ is a probability measure satisfying $\mu(\bm{x})=\sum_{\bm{y}\in\bm{\mathcal{S}}}\text{Pr}_{\bm{x},\bm{y}}\mu(\bm{y})$. A realization of a Markov chain lies in a space constructed by a countable product $\Pi_{i=0}^{\infty}\bm{\mathcal{S}}_{i}$, where each $\bm{\mathcal{S}}_{i}$ is a copy of $\bm{\mathcal{S}}$ equipped with a copy of the $\sigma$-algebra of $\bm{\mathcal{S}}$ [Meyn:1253345]. Intuitively, a realization of a Markov chain is a set of values of $\mathbf{X}_{0},\mathbf{X}_{1},\ldots$, which are generated based on the transition matrix $\mathbf{P}$. We say that an event $A$ happens almost surely if it happens with probability one, i.e., $\mathbb{P}(A)=1$. This well-known definition is motivated by events $A$ whose complements $A^{c}$ are non-empty, i.e., $A^{c}\neq\emptyset$, but have a zero measure. A sequence of probability measures $\langle{\mu_{k}}\rangle_{k\in\mathbb{Z}_{\geq 0}}$ defined on $\bm{\mathcal{S}}$ converges weakly to a probability measure $\mu$ if for any bounded continuous function ${f}:\bm{\mathcal{S}}\to\mathbb{R}$, we have $\int fd\mu_{k}\xrightarrow{k\to\infty}\int fd\mu$ [klenke2013probability]. A sequence $\langle\epsilon_{k}\rangle_{k=0}^{\infty}$ of positive constants that converges to zero is called vanishing.
@@
[roth2013stochastic, Theorem 3.5 and Corollary 3.9] For a vanishing sequence of $\epsilon>0$, let $\langle\langle\mathbf{X}^{\epsilon}_{k}\rangle{{}_{k=0}^{\infty}}\rangle_{\epsilon>0}$ be GSAPs for a good upper semicontinuous differential inclusion $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$. Assume that for each $\epsilon$, $\langle\mathbf{X}^{\epsilon}_{k}\rangle_{k=0}^{\infty}$ is a Markov chain and let ${\mu}^{\epsilon}$ be an invariant probability measure of $\langle\mathbf{X}^{\epsilon}_{k}\rangle{{}_{k=0}^{\infty}}$. Let $\mu$} be a limit point of $\langle{\mu}^{\epsilon}\rangle_{\epsilon>0}$ in the topology of weak convergence. Then the support of $\mu$} is contained in $\mathrm{BC}(\bm{\Phi})$, where $\bm{\Phi}$ is the dynamical system induced by $\dot{\bm{x}}\in\bm{\mathcal{V}}(\bm{x})$.
@@ VII-A Proof of Theorem 1
Assume by contradiction that $\lim_{k\to\infty}{{\mu}}^{\epsilon_{k}}(\bm{\mathcal{{O}}})\not\to 1$. This implies that for some positive constant $\delta<1$, there exists a subsequence $\langle\alpha_{k}\rangle_{k=1}^{\infty}=\alpha_{1},\alpha_{2},\ldots$ of $\langle\epsilon_{k}\rangle_{k=0}^{\infty}$ such that ${{\mu}}^{\alpha_{k}}(\bm{\mathcal{{O}}})<\delta$ for each $k$. As the state space $\bm{\mathcal{X}}_{0}$ is compact, the family of probability measures defined on the space $\bm{\mathcal{X}}_{0}$ is tight [klenke2013probability]. Then, in view of Prokhorov’s theorem, every sequence of the probability measures, including $\langle{\mu}^{\alpha}\rangle$ has a weak limit point, which is a probability measure $\mu^{\prime}$}. This implies that there exists a subsequence $\langle\beta_{k}\rangle_{k=1}^{\infty}=\beta_{1},\beta_{2},\ldots$ of $\langle\alpha_{k}\rangle_{k=1}^{\infty}$ such that $\lim\inf_{k\to\infty}{{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})\geq{{\mu^{\prime}}}(\bm{\mathcal{{O}}})$. On the other hand, in view of A1, the support of every limit point is contained in the Birkhoff center $\mathrm{BC}(\bm{\Phi})$, i.e., ${\mu}^{\prime}(\bm{\mathcal{O}})=1$. This means that there exists some $k^{*}$ such that for all $k>k^{*}$ ${{\mu}}^{\beta_{k}}(\bm{\mathcal{{O}}})>\delta$, a contradiction. ∎
@@ VII-C Proof of Lemma 1
Under 1, let $d$ denote the minimum distance between thresholds, i.e., $d=\min_{i,j\in[\mathsf{p}+\mathsf{p}^{\prime}]}|\tau_{i}-\tau_{j}|.$ Then, for any population size $\mathsf{N}$ satisfying $\mathsf{N}d>1$ and any $\bm{x}\in\bm{\mathcal{X}}_{ss}$, the sum $\bm{1}^{\top}\bm{x}$ belongs to at most one interval of the form $\tau_{i},\tau_{i}+\frac{1}{\mathsf{N}})$ with $i\in[\mathsf{p}+\mathsf{p}^{\prime}]$.
@@
The proof of Lemma 1 is inspired by the steps taken in [roth2013stochastic, Example 4.1].
@@
By considering the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence, it suffices to show that there exist $\langle\bm{\mathcal{V}}^{\frac{1}{\mathsf{N}}}\rangle_{{\frac{1}{\mathsf{N}}}}$ and $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ such that the conditions in Definition 4 are met. The first condition of Definition 4 can be easily verified. Let $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})$ denote the expected increment per time unit of the Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ when the Markov state is at $\bm{x}^{\mathsf{N}}$. Since there are $\mathsf{N}$ transitions per unit time, we have $\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x}^{\mathsf{N}})=\smash{\mathsf{N}\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}]}$. The $p^{\text{th}}$ element of $\bm{\nu}^{\frac{1}{\mathsf{N}}}$, $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, which is the expected increment associated with the subpopulation $p$, equals the sum of the multiplication of each possible change with its probability. In view of (6), there are two possible changes, i.e., $\frac{1}{\mathsf{N}}$ step increment or $\frac{1}{\mathsf{N}}$ step reduction, which results in ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=\mathsf{N}\big(\frac{1}{\mathsf{N}}(\rho_{p}-x^{\mathsf{N}}_{p})(2-s^{*}(p,\bm{x}^{\mathsf{N}},2))-\frac{1}{\mathsf{N}}(x^{\mathsf{N}}_{p}(s^{*}(p,\bm{x}^{\mathsf{N}},1)-1))\big)$ and consequently,
${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}}^{\mathsf{N}})=$ $\rho_{p}(2-s^{*}(p,\bm{x}^{\mathsf{N}},2))-x^{\mathsf{N}}_{p}$
$+x^{\mathsf{N}}_{p}\left(s^{*}(p,\bm{x}^{\mathsf{N}},2)-s^{*}(p,\bm{x}^{\mathsf{N}},1)\right).$ (18)
The second condition of Definition 4 is satisfied by $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}=$ $\smash{\mathsf{N}(\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-\mathbb{E}[\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\mid\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}=\bm{x}^{\mathsf{N}}])}$ resulting in $\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k+1}-\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}-{\frac{1}{\mathsf{N}}}\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k+1}={\frac{1}{\mathsf{N}}}\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$. Now, we investigate Condition 3. To show that Condition 3 is satisfied, for a given $\delta$, we need to provide an $\epsilon_{0}$ satisfying eq. 5. We propose an $\epsilon_{0}$ which satisfies $\epsilon_{0}<\min\{d,\delta\}.$ Then, given A1, for any $\bm{x}\in\bm{\mathcal{X}}_{ss}$ there exists at most one threshold, say $\tau_{l}$, $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$, that satisfies $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0})$, with $\frac{1}{\mathsf{N}}\leq\epsilon_{0}.$ In what follows, we show that Condition 3 holds for all $\bm{x}\in\bm{\mathcal{X}}_{ss}$ that satisfy $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$; the other values of $\bm{x}$ can be handled similarly. Let $\varepsilon_{x}=\bm{1}^{\top}\bm{x}-\tau_{l}$. Then given $\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$, we have $\varepsilon_{x}<\epsilon_{0}$ and, in turn, $\varepsilon_{x}<\delta.$ Consider the point $\bm{y}$ whose all elements, except some element $j$, are equal to the corresponding elements of $\bm{x}$, and the $j^{\text{th}}$ element, $y_{j}$, equals $x_{j}-\varepsilon_{x}$ (the value of $x_{j}$ should be greater than or equal to $\varepsilon_{x}$). The term $|\bm{x}-\bm{y}|$ then equals $\varepsilon_{x}$, and hence the point $\bm{y}$, which satisfies $\bm{1}^{\top}\bm{y}=\tau_{l}$, belongs to the $\delta$-neighborhood of $\bm{x}$, i.e., $|\bm{x}-\bm{y}|<\delta$. For all $p\neq l$, we then have $\mathcal{V}_{p}({\bm{y}})=\rho_{p}(2-s^{*}(p,{\bm{y}},2))-{y}_{p}$ and for $p=l$, $\mathcal{V}_{p}=[-{y}_{p},\rho_{p}-{y}_{p}]$. On the other hand, in view of (18) and $\exists l\in[\mathsf{p}+\mathsf{p}^{\prime}](\bm{1}^{\top}\bm{x}\in[\tau_{l},\tau_{l}+\epsilon_{0}))$, the term ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ for $p=l$ equals $\rho_{l}-2x^{\frac{1}{\mathsf{N}}}_{l}$ (resp. 0) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation. For $p\neq l$, ${\nu}^{\frac{1}{\mathsf{N}}}_{p}({\bm{x}})$ equals $\rho_{p}(2-s^{*}(p,\bm{x},2))-x_{p}$, and it can be easily shown that it is also equal to the member of the singleton $\mathcal{V}_{p}(\bm{x})$ defined in (7). By $\bm{z}=\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})$ and choosing $\bm{v}$ to be a member of $\bm{\mathcal{V}}({\bm{y}})$ with $v_{l}({\bm{y}})=\rho_{l}-2{y}_{l}$ (resp. $v_{l}({\bm{y}})=0$) if $l$ corresponds to a coordinating (resp. an anticoordinating) subpopulation, the term $|\bm{z}-\bm{v}|$ equals $|\bm{v}({\bm{y}})-\bm{\nu}^{\frac{1}{\mathsf{N}}}(\bm{x})|=|{\bm{y}}-\bm{x}|$ which is less than $\delta$. This implies that the relation $\inf_{\bm{v}\in\bm{\mathcal{V}}(\bm{y})}\!|\bm{z}-\bm{v}|<\delta$ holds and, in turn, Condition 3 is satisfied. As for the last condition, in view of Example 4.1 in [roth2013stochastic], it can be shown that $\langle\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ is a martingale difference sequence. On the other hand, for all $k$, $|\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}|\!\leq\!\sqrt{\Sigma_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}(1+\rho_{l})^{2}}$ and hence $\mathbf{U}^{\frac{1}{\mathsf{N}}}_{k}\!$ is uniformly bounded. Therefore, the conditions asserted in [roth2013stochastic, Proposition 2.3] are satisfied, and, consequently, the last condition holds. ∎